OpenStudy (anonymous):

a to-gun pilot, practising radar avoidance maneuvers, is manually flying horizontally at 1300km/h, just 35m above the level ground. suddenly, the plane encounters terrain that slopes gently upward at 4.3 degree. an amount difficult to detect visually . how much times does the pilot have to make a correction to avoid flying into ground? how to do this problem...???

OpenStudy (anonymous):

its top-gun pilot.... typing mistake...

OpenStudy (anonymous):

imagine that the slope rising at angle 4.3 is the hypotenuse of the triangle ABC, then the height of the triangle AC is 35 (because that is the altitude of the jet above the ground), then the length from the point of rising till the point where the jet would crash BC is: \[tan\theta= AC/BC\] \[tan(4.3)= 35/BC\] \[BC= 35/tan(4.3) = 465.485 m\] So the distance to the point where the jet might collide is 465.485 m. you know the speed of the jet, 1300 km/h, that is 361.111 m/s, dividing the length BC by speed: \[465.485/361.111 = 1.289 \space seconds\] very interesting problem, a pilot must be good in flying :)

OpenStudy (anonymous):

thank you onaogh